2. For a line segment ST, one endpoint is S(-1/2), 4) and the midpoint is
M(6, -3/4).
Find the coordinates of the other endpoint T.
Thank you very much!!!! Answer by edjones(8007) 
 For x: 6-(-1/2)=6 1/2
6+6 1/2=12 1/2
.
For y: (-3/4)-4=-4 3/4
-3/4+(-4 3/4)=-5 1/2
.
(12 1/2), (-5 1/2) other end point.
.
Ed